OpenStudy (anonymous):

Evaluate: 20 / 2y - y, for y = 5

OpenStudy (amistre64):

20/2(5) -5 = 20/10 - 5 = 10-5 = 5

OpenStudy (amistre64):

\begin{align} 20/2(5) - (5)&=20/10 - 5\\ &=10-5\\ &=5 \end{align} maybe

OpenStudy (anonymous):

Do I do anything with that stupid y?

OpenStudy (amistre64):

which y is the stupid one? :)

OpenStudy (anonymous):

Haha, the 2y. Or is this problem saying that the 2y is actually a hidden 5, and it is then subtracting another 5?

OpenStudy (amistre64):

the y itself is just standing in line holding someone place tell they get back; it just so happens that its holding the 5s place :)

OpenStudy (anonymous):

That is very clever! :) Thanks again, you're awesome!

OpenStudy (amistre64):

yep :)

OpenStudy (anonymous):

Hello guys, can you help me with an algebra problem? -3y^3-5y-6(y+2y^3)

OpenStudy (amistre64):

the best place to get your question noticed is if you post it up on the left side in the "ask a question" box.